OpenStudy (anonymous):

how do i solve f+4 and under is g and it =6 for f

OpenStudy (anonymous):

\[\frac{f+4}{g}=6\]?

OpenStudy (anonymous):

yup how would i solve

OpenStudy (anonymous):

what is g? is there any infomation?

OpenStudy (anonymous):

system of equation. so you have 6g = f+4 so g = f+4/6

OpenStudy (anonymous):

That is my own idea then you can plug it in G,

OpenStudy (anonymous):

i want to see full question

OpenStudy (anonymous):

multiply by g f+4=6g subtract 4 f=6g-4

OpenStudy (anonymous):

f+4 and under is g =6

OpenStudy (anonymous):

so that would be th e answer mukushla

OpenStudy (anonymous):

i don't know why your teacher gives this question it's meaningless

OpenStudy (anonymous):

she said so herself she said its a stupid lesson

OpenStudy (anonymous):

answer is f=6g-4 but do u know how to get it?

OpenStudy (anonymous):

yeas

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

maybe yulitzi is not college student..